OpenStudy (anonymous):

Calculate the values of x if @ganeshie8

OpenStudy (anonymous):

\[3^{2x+1}-3^{x+2}=3^{x+3}-3^4\]

OpenStudy (anonymous):

let \(t=3^x\) and make a quadratic in terms of \(t\)

OpenStudy (anonymous):

\[t(3^3)-t(3^2)-t(3^3)=-3^4\]

OpenStudy (anonymous):

\[t(3^3-3^2-3^3)=-3^4\]\[t(9)=81\]

OpenStudy (anonymous):

\[t=9\]

OpenStudy (anonymous):

\[x=2\]

OpenStudy (anonymous):

Like this? @mukushla

OpenStudy (anonymous):

except that\[3^{2x+1}=3t^2\]

OpenStudy (anonymous):

\[3t^2-9t-27t+81=0\]

OpenStudy (anonymous):

\[3t^2-36t+81=0\]

OpenStudy (anonymous):

\[3(t^2-12t+27)=0\]

OpenStudy (anonymous):

\[t=3,t=9\]

OpenStudy (anonymous):

\[3^x=3^1,3^x=3^2\]

OpenStudy (anonymous):

\[x=1,x=2\]

OpenStudy (anonymous):

Thnx @mukushla

OpenStudy (anonymous):

no problem